The length of a house - and - lot for sale is 6 m more than its width. The floor area is surrounded by a 3 - meter walkway. The total area of the walkway is 288 m^2. What are the dimensions of the lot?

Answer by ankor@dixie-net.com(22740)

The length of a house - and - lot for sale is 6 m more than its width.
Assuming the length of the house is 6 meters longer than the width
L = w + 6
:
The floor area is surrounded by a 3 - meter walkway.
Over all dimensions
(L+6) by (w+6)
:
The total area of the walkway is 288 m^2.
Overall area - house area = walkway area
(L+6)(w+6) - Lw = 288
Replace L with (w+6)
(w+12)*(w+6) - w(w+6) = 288
FOIL
w^2 + 6w + 12w + 72 - w^2 - 6w = 288
combine like terms
w^2 - w^2 + 18w - 6w = 288 - 72
12w = 216
w = 216/12
w = 18 m is the width of the house
then
18+6 = 24 m is the length of the house
therefore
24+6 = 30 m is the lot length
and
18+6 = 24 m is the lot width
:
:
Check
lot - house = walkway
(30*24) - (24*18) =
720 - 432 = 288 sq meters